fred is analyzing the cost of producing two different items at an electronics company. an electrical sensing device uses 5 grams of copper and requires 3 hours to assemble. a smaller sensing device made by the same company uses 4 grams of copper but requires 5 hours to assemble. the first device has a production cost of $27. the second device has a production cost of $32. how much does it cost the company for a gram of this type of copper? what is the hourly labor cost at this company? (assuming that production cost is obtained by adding the copper cost and the labor cost)
Answer by stanbon(75887)

fred is analyzing the cost of producing two different items at an electronics company.
An electrical sensing device uses 5 grams of copper and requires 3 hours to assemble.
A smaller sensing device made by the same company uses 4 grams of copper but requires 5 hours to assemble.
======================================
The first device has a production cost of $27.
The second device has a production cost of $32.
======================================
How much does it cost the company for a gram of this type of copper?
What is the hourly labor cost at this company? (assuming that production cost is obtained by adding the copper cost and the labor cost)
-----
5c + 3a = 27
4c + 5a = 32
--------
Modify:
25c + 15a = 5*27
12c + 15a = 3*32
-------------------
Solve for "c"
13c = 39
c = 3 dollars (cost of one gram of copper)
=================
Cheers,
Stan H.
=================